. (4) Solve the system of equations.
6x – 2y = 2
9x – 3y = 1

Answer by venugopalramana(3286)

6x – 2y = 2....................I
9x – 3y = 1..................II
3*EQN.I-2*EN.II
18X-6Y-18X+6Y=6-2=4
0=4
HENCE THE EQNS.ARE DEPENDENT AND INCONSISTENT.HENCE THERE IS NO SOLUTION